If a train travels 40 miles per hour from one city to another, it reaches 10 minutes late and if it travels 30 miles per hour then it reaches 16 minutes late. What is distance between two cities.
Answer by mananth(16946)

let t be the exact time it has to reach
x/40 =t+1/6
x/40-1/6= t
x/30 =t+4/15
x/30 -4/15=t
x/40-1/6 = x/30-4/15
x/40-x/30 = -4/15+1/6
(3x-4x)/120 = (-8+5)/30
-x= 4(-3)
-x=-12
x=12 miles